OpenStudy (anonymous):

i need help with this math question let f(x) = x+4/x+1, x is not equal -1, and g (x)=x-2/x-4, x is not equal to 4. find the set of values of x such that f(x) is equal to or less than G(x)

OpenStudy (freckles):

so it is asking you to solve f(x)<=G(x)

OpenStudy (freckles):

but I don't think G(x) has been given... Is g(x)=G(x)?

OpenStudy (freckles):

or is there another type of relationship between g and G

OpenStudy (anonymous):

g(x)is the same as G(x) sory just a typo

OpenStudy (freckles):

so you need to solve f(x)<=g(x)

OpenStudy (freckles):

that is \[\frac{x+4}{x+1} \le \frac{x-2}{x-4}\]

OpenStudy (freckles):

First step: Make one side of the inequality 0 by moving one fraction to the other side by either addition/subtraction which ever one works in this case

OpenStudy (anonymous):

alright and second step

OpenStudy (freckles):

combine the fractions

OpenStudy (freckles):

find the zeros of the fraction and where the fraction doesn't exist

OpenStudy (freckles):

draw a number line and test the intervals around each of those numbers

OpenStudy (freckles):

so first what do you get as the inequality after the first step

OpenStudy (anonymous):

im stuck on that

OpenStudy (freckles):

\[f(x) \le g(x) \\ \text{ subtract } g(x) \text{ on both sides } f(x) -g(x) \le 0\]

OpenStudy (freckles):

now combine those fractions

OpenStudy (anonymous):

2/-3

OpenStudy (freckles):

him you seem to be missing the variables...

OpenStudy (freckles):

you have \[\frac{x+4}{x+1}-\frac{x-2}{x-4} \le 0\]

OpenStudy (freckles):

try combining the fractions on the left

OpenStudy (anonymous):

is it x-14/(x-4)(x+1)